OpenStudy (anonymous):

Another logarithm. log[5](x+116)=4-log[5](x-4)

OpenStudy (anonymous):

Move all your logs to the same side, and combine them by the property that: \[log_b(a) + log_b(c) = log_b(a\cdot c))\]

OpenStudy (anonymous):

I'm confused by the negative in front of the second log

OpenStudy (anonymous):

Just add \(log_5(x-4)\) to both sides of the equal sign.

OpenStudy (anonymous):

That will zero out the one on the right, and make a new, positive one on the left you can combine.

OpenStudy (anonymous):

log[5](x+116/x-4)=4 and solve from there??

OpenStudy (anonymous):

Not divide, multiply.

OpenStudy (anonymous):

\[log_5((x+116)(x-4)) = 4\]

OpenStudy (anonymous):

(x+116)(x-4)=5^4 and go from there?